How can I show the equation use for the solution?
A coffee merchant has coffee beans that sell for $9 per
pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?

Answer by mathispowerful(115)


The idea is that the total worth of coffee will remain unchanged before and after mixture.
Let x represent the amount of $9-coffee
y represent the amount of $12-coffee
then
x + y = 100
9x + 12y = 11.25 * 100 = 1125
so we have
x + y = 100
9x + 12y = 1125
multiply -9 on both sides of first equation
-9x - 9y = -900
add to second equation
3y = 225
y = 75 lbs
and x = 100 -75 = 25 lbs
so we need to mix 25 lbs of $9-coffee with 75 lbs of $12-coffee.